Circular No. 8756

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 OF_{2} (BROUGHTON) An object reported as asteroidal when discovered by J. Broughton

(Reedy Creek, Queensland, Australia, 0.25-m reflector; discovery observation tabulated below from MPEC 2006-O13 and MPC 174866), and thus given the minor-planet designation 2006 OF_{2}, has been found by C. W. Hergenrother to be cometary on CCD frames taken with the 1.54-m reflector at Catalina, whose 60-s R-band images on Sept. 20.11 UT show the comet to possess a small, condensed coma 7′′ in diameter and magnitude 18.2 with no tail. Additional images taken by Hergenrother in good seeing (< 1′′) on Sept. 26.16 show an obvious condensed asymmetric coma of diameter 7′′ that is elongated towards the northeast.

2006 UT  α_{2000}  δ_{2000}  Mag.
July 17.65728  20^{h}43^{m}51.^{s}49  −27^{o}37′11′′.2  18.2

Additional astrometry, the following hyperbolic orbital elements, and an ephemeris appear on MPEC 2006-S91.

Epoch = 2008 Sept. 11.0 TT

T = 2008 Sept.15.5510 TT  ω = 95.^{o}6032
e = 1.001046  Ω = 318.5093 }2000.0
q = 2.431387 AU  i = 30.1670

1999 RT_{214}

K. S. Noll, Space Telescope Science Institute; W. M. Grundy, Lowell Observatory; H. F. Levison, Southwest Research Institute; and D. C. Stephens, Johns Hopkins University, report the detection of a binary companion to the transneptunian object 1999 RT_{214} (cf. MPEC 1999-X27, MPO 50421), which is a member of the low-inclination classical TNO population. The observations were made during 2006 July 25.233–25.309 UT with the High Resolution Camera of the Advanced Camera for Surveys on the Hubble Space Telescope (HST), using the clear filters with one 300-s exposure at each of four dithered positions on the detector. The two components are clearly resolved in each image and in the coadded image, separated by an angular distance of 0′′.115 ± 0′′.001 and differ in brightness by 0.8 mag. The fainter component lies at a position angle of 89^{o}.4 ± 1^{o}.3 from the primary. The projected separation of the objects in the sky plane is 3310 ± 30 km. The HST corrected for parallax and tracked both components of 1999 RT_{214} as they moved together at an average rate of 0′′.045/min.

2006 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green